I have tried to solve this problem. It is dealing with distributive properties. I need some help. I would like to see the step by step solution for this problem so I can see where I made my mistake. Here is the problem.
Solve for x: 3(4x + 1)-2=9x + 1
Thanks for your help

Answer by nerdybill(2448)

3(4x + 1)-2=9x + 1
First, distribute the 3 to terms inside the parenthesis:
12x + 3 -2 = 9x + 1
Combine like-terms on the left:
12x + 1 = 9x + 1
Subtract 9x from both sides:
3x + 1 = 1
Subtract 1 from both sides:
3x = 0
Finally, divide both sides by 3:
x = 0